A bag contains 4 red, 3 pink, and 6 green balls. Two balls are drawn but the first ball drawn is not replaced.





a. find P(red, then pink) b. find P(pink then pink)

1
Expert's answer
2022-05-02T02:51:27-0400

a. Initially there are 4 + 3 + 6 = 13 balls in the bag, 4 of them are red.

The probability of choosing a red ball on the first draw is

"P_1=\cfrac{4}{13}."

Since we do not put it back, there are 12 balls, 3 of which are pink. The second pink has probability

"P_2=\cfrac{3}{12}."

So, the probability that we draw first red and second pink, is

"P=P_1\cdot P_2=\cfrac{4}{13}\cdot\cfrac{3}{12}=\cfrac{1}{13}."


b. The probability of choosing a pink ball on the first draw is

"P_1=\cfrac{3}{13}."

Now there are 12 balls, 2 of which are pink. The second pink has probability

"P_2=\cfrac{2}{12}."

The probability that we draw first pink and second pink, is

"P=P_1\cdot P_2=\cfrac{3}{13}\cdot\cfrac{2}{12}=\cfrac{1}{26}."